# Lattice Boltzmann methods for soft flowing matter

### v1

@@ V.1 Soft flowing crystals
We initially consider the case of fluid droplets flowing in a microfluidic channel, a system simulated using a color-gradient LB incorporating near-contact interactions, as described in section III.3.1 Montessori et al. (2019b, 2020). Such droplets are usually produced in flow-focusing devices through an emulsification process, in which their generation follows from the periodic pinch-off of the liquid jet of the dispersed phase by the stream of the continuous phase through a small orifice separating the inlet and the outlet chambers Marmottant and Raven (2009); Raven and Marmottant (2009); Garstecki et al. (2004). This emulsion is stabilized by dispersing surfactant molecules preventing coalescence. A sketch of the geometry used in our simulations and the droplet arrangement observed in the channel are shown in Fig.3. These devices are particularly suited for the design of highly monodisperse droplets which are of relevance, for example, for the assembly of droplet-based soft templates with a well-defined structure Nan et al. (2024); Xu et al. (2005); Costantini et al. (2014).
@@
Typical examples of ordered soft materials are soft flowing crystal (also termed as microfluidic crystals Marmottant and Raven (2009)), where monodisperse fluid droplets added in a self-repeating way are found to self-assemble in regular patterns, closely resembling the hexagonal order of solids. In simulations, droplet formation can be controlled by tuning the ratio of dispersed-to-continuous inlet flow, defined as $\chi=u_{d}/2u_{c}$ (where $u_{d}$ and $u_{c}$ are the speed of dispersed and continuous phases) and the dimensionless near-contact number ${\cal N}_{c}=A\Delta x/\sigma$ (see also Eq.42). Here, $A$ is a constant setting the magnitude of the repulsive force and $\Delta x$ is the minimum distance between two interfaces in close contact, generally ranging from $10$ to $100$ nanometers. If ${\cal N}_{c}\ll 1$ droplets coalesce, otherwise repulsive near-contact forces prevail and droplet merging is inhibited.
@@
In Fig.13a-d we show a variety of crystal-like structures confined in microfluidic channels obtained from LB simulations and experiments of air bubbles in water Marmottant and Raven (2009); Raven and Marmottant (2009). Despite the difference in density ratio, the two systems exhibit very similar features. In particular, the emulsion displays a number of phases comprising a three-row configuration (hex-three, (a)), where a regular array of droplets accommodates along three parallel rows, two two-row ones (wet (b) and dry (c) hex-two) and a single-row structure (hex-one, (d)), in which highly-packed droplets arrange in a neat line. Besides being capable of describing the formation and ordering properties of fluid emulsions under confinement, LB simulations also capture the behavior of the flow rate curve and, remarkably, the non-trivial transition from the hex-two to the hex-one regime (see Fig.13e,f). Indeed, experiments show that, for a gas-liquid foam, $Q_{g}\propto P_{g}^{\alpha}$, where $Q_{g}$ and $P_{g}$ are flow rate and pressure of the gas, and $\alpha=3/2$ for small values of $P_{g}$ Marmottant and Raven (2009). On the contrary, for higher values of $P_{g}$, $Q_{g}$ displays a sharp decrease due to the augmented friction of the interfaces with the walls, an effect signaling the transition to the hex-one phase (Fig.13e). These results are reproduced, with high accuracy, by LB simulations of fluid emulsions, where the flow rate $Q_{d}$ of the dispersed phase follows a power law of $\phi$ (which plays the same role as the gas pressure) with an almost equal value of the exponent. In agreement with experiments, simulations also show that the transition from the hex-two to the hex-one phase is marked by a T1 event Graner et al. (2008); Marmottant et al. (2008); Weaire and Hutzler (2001) (corresponding to a topological rearrangement of neighboring droplets) and a transition front, separating the two phases and comoving with the flowing emulsion.
@@
Figure 13: (a)-(d) Simulated (top rows) and experimental (bottom rows) structures of flowing droplets in a microfluidic channel for $\chi=1$ (a) (hex-three), $\chi=1.5$ (b) (wet hex-two), $\chi=2$ (c) (dry hex-two) and $\chi=3.6$ (d) (hex-one). Here $\chi=u_{d}/2u_{c}$, being $u_{d}$ and $u_{c}$ the speed of dispersed and continuous phases. (e)-(f) Flow curves of experimental Marmottant and Raven (2009) and simulated Montessori et al. (2020) soft flowing crystals. In (e) $P_{g}$ and $Q_{g}$ are the gas pressure and gas flow rate. For values of $P_{g}$ lower than $4$kPa, the mixture is in the hex-two regime, where the flow rate scales as $Q_{g}\propto P^{1/\alpha}$ and $\alpha\simeq 2/3$. At $P_{g}\simeq 4$kPa, $Q_{g}$ displays a sharp transition towards the hex-one regime. In (f) $Q_{d}=\nu V_{d}$ is the flow rate of the dispersed phase, where $\nu$ is the frequency of droplet generation and $V_{d}$ is the volume of the generated droplet. The flow rate follows a power law for low values of $\chi$, while for $\chi\simeq 3.5$ a decrease of $Q_{d}$ marks the transition from the hex-two to the hex-one phase. The transition front is indicated by a white dotted line in the top inset. The figures are reproduced with permission from Ref.Montessori et al. (2019b, 2020).
@@
LB simulations of microflows have also been capable of describing morphology and assembly of droplets flowing in convergent and divergent channels Montessori et al. (2019b); Gai et al. (2022); Y. et al. (2016); Luo et al. (2021); Montessori et al. (2021a, b), as the ones shown in Fig.14. More specifically, numerical results suggest that the degree of monodispersity and the arrangement of a confined dense emulsion can be controlled by properly tuning the opening angle ($\alpha$) of a divergent channel and Capillary number ($Ca$) (see Fig.14). Indeed, while for $\alpha\lesssim 40^{\circ}$ the emulsion is generally monodisperse (Fig.14a) regardless of the values of $Ca$, for $45^{\circ}\lesssim\alpha\lesssim 60^{\circ}$ the generated emulsion becomes bidisperse, an effect amplified at high values of $Ca$ (see Fig.14b,c) because breakup events are more likely to occur. The mechanics leading to droplet rupture is highlighted in Fig.14e-n. The orange droplet gradually stretches because of the combined effect of i) the normal force of the incoming droplet (yellow), ii) the extensional force of the channel expansion, and iii) the downstream droplet (red) acting as a ”wall”. Once a critical elongation is overcome, the orange droplet fragments produce two smaller droplets moving towards the upper and lower part of the channel. Then the process self-repeats; the yellow droplet fills the gap left by the orange one and acts as a ”wall” droplet, while a newly generated one will stretch and finally break following a protocol similar to the orange one. Interestingly, further increasing $\alpha$ restores a monodisperse emulsion, once again regardless of the values of $Ca$ (see Fig.14d). This counter-intuitive behavior is due to the fast recovery of the circular shape in the main channel because of larger aperture angles. Such an expansion favors the transversal displacement of the incoming droplet, thus suppressing the squeezing effect. Note finally that, for all geometries considered, increasing $Ca$ generally leads to the formation of a foamy-like emulsion; this is because, in this regime, morphological deformations become easier and, once breakups occur, the resulting smaller droplets fill the voids among neighboring ones.
@@
Figure 14: (a)-(d). Droplet assembly in diverging channels of opening angle $\alpha$ equal to $22.5^{\circ}$ (a), $45^{\circ}$ (b), $60^{\circ}$ (c), $90^{\circ}$ (d) and three different values of Ca. The histograms show the distribution of droplet diameters $D_{e}$ for $Ca=0.04$ (dashed line), $Ca=0.1$ (dotted line) $Ca=0.16$ (full line). Note that, for $45^{\circ}\lesssim\alpha\lesssim 60^{\circ}$, the emulsion displays a bidisperse structure (disordered), while for values out of that range, the emulsion is essentially monodisperse (ordered). (e)-(n). Pinch-off mechanism from LB simulations (e-h) and experiments (i-n) at $Ca\simeq 0.08$ and $\alpha=45^{\circ}$. The split of the orange droplet is caused by the confining effect of the yellow droplet (the ”pincher”) and the red one (the ”wall”). The figure is reproduced with permission from Ref.Montessori et al. (2021b).
@@ V.2 Soft granular media
Figure 15: Left panel, (a)-(c). Phase diagram of a soft granular medium, where $N$ is the number of cores and $\phi_{d}$ is the volume fraction they occupy. Insets (a), (b) and (c) represent three possible configurations, where (a) indicates weakly deformed shells, (b) highly deformed ones, and (c) a linear chain. Right panel, (d)-(e). Experimental realization of a soft granular medium (d) formulated using three Newtonian fluids. A result obtained from LB simulations (e) using three-component fluid, i.e. the dispersed phase (white droplets), the lubricating phase (blue), and an external fluid (black). Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license and https://creativecommons.org/licenses/by/3.0/ CC 3.0 license respectively. No changes were made from the original figure from Ref.Montessori et al. (2021c); Guzowski and Garstecki (2015).
Recent experiments have shown that these materials exhibit a complex phase diagram, where multiple topologies can be obtained by varying the volume fraction $\phi_{d}$ and the number $N$ of the internal droplets (Fig.15a-c). Indeed, one can identify a region where the packing mildly deforms the emulsion (for $\phi\simeq 0.5$) alongside a wider phase where the emulsion shows a significant departure from the spherical shape (for $0.6\lesssim\phi\lesssim 0.8$). For higher values of $\phi_{d}$, the internal droplets (also termed as cores) arrange into a stable linear chain encapsulated within an elongated shell. Such a picture decisively depends on the grain structure as well as on the deformability of the cores, features controlling the formation of highly packed structures with a volume fraction typically well above the close packing limit of hard spheres.
@@
Figure 16: (a) Dynamic patterns observed in a soft granular medium by varying $Q_{c}/Q_{d}$. The medium consists of water-in-oil emulsion surrounded by a further immiscible phase (fluorinated fluid). The left column shows simulation results, the central one shows the corresponding experiments and the right one refers to a water (blue)-oil (transparent) mixture. The color bar refers to the fluid velocity in simulations. (b) Mechanism leading to the jet breakup. If a doublet enters the narrowing and unfolds, the droplet chain is stabilized. Otherwise, high velocity gradients in the surrounding of the doubled (see the pink arrow) destabilize the chain. Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license. No changes were made from the original figure from Ref.Bogdan et al. (2022).
LB simulations prove, once again, instrumental to capture crucial aspects of the physics of such systems. In Fig.16a, for example, a series of experiments about the formation of dynamic modes of a confined granular medium (a densely packed monodisperse emulsion focused by an external immiscible phase through an orifice) are presented, while the numerical simulations are performed using the color gradient LB described in section III.3.1. Upon varying the flow ratio $Q_{c}/Q_{d}$ between the continuous and the dispersed phases, both experiments and simulations show that one can distinguish four modes, consisting of (i) a thick, weakly oscillating jetting phase, (ii) a thin, strongly oscillating jet with occasional breakups, (iii) an irregular dripping phase of polydisperse double emulsions, and (iv) a regular dripping phase of monodisperse double emulsions. In stark contrast, Newtonian liquids (such as water in oil mixture) only show two primary modes, i.e. dripping and jetting, although other regimes have been reported Utada et al. (2007).
Besides describing at a high level of accuracy (an almost one-to-one mapping to experiments) these complex structures, LB simulations also unveil the mechanism causing the jet breakup and leading to the dripping phase, as shown in Fig.16b. This is essentially based on the droplet rearrangement upon entering the constriction and on the magnitude of the associated velocity gradients. Indeed, when a doublet (i.e. a pair of droplets) enters the narrowing, the surrounding continuous phase accelerates to conserve the flux. If the doublet unfolds into a chain, the velocity gradients remain weak and the chain is stabilized; on the contrary, if the doublet survives, increasing velocity gradients around it destabilize the jet leading to the breakup.

### v2

@@ V.1 Soft flowing crystals
We initially consider the case of fluid droplets flowing in a microfluidic channel, a system simulated using a color-gradient LB incorporating near-contact interactions, as described in section III.3.1 Montessori et al. (2019b, 2020). Such droplets are usually produced through an emulsification process, in which their generation follows from the periodic pinch-off of the liquid jet of the dispersed phaseby the stream of the continuous phase. Typical microfluidic platforms used for droplet generation are T-junctions, flow focusers and divergent channels.
T-junction.
As benchmark test, we start from a T-junction device where a droplet (blue region) is immersed in a continuous phase (yellow region) and flows within the microchannel (see Fig.14a). The two components have equal density while the viscosity of the continuous phase is approximately five times larger than that of the dispersed phase, in order to match experimental data of Cubaud and Mason (2008). By changing the capillary number $Ca_{1}$ and $Ca_{2}$ of both phases, one can distinguish four flow regimes: a dripping regime, consisting of elongated droplets, for $Ca_{1}<0.1$ and $Ca_{2}<0.1$, where capillary forces dominate over inertial ones (top row in Fig.14a and $C$-zone in Fig.14b); two jetting regimes for $10.1<Ca_{1}<$ 1 and $10^{-3}<Ca_{2}<10^{-2}$, where low viscous forces of the continuous phase causes the formation of a thin jet of fluid delaying the droplet generation (second and third rows in Fig.14, and $B$-zone in Fig.14b); a tubing regime for $Ca_{1}\simeq 10^{-1}$ and $Ca_{2}<10^{-2}$, where a jet of the dispersed phase flows in parallel with the continuous one in the microchannel (fourth row in Fig.14a and $D$-zone in Fig.14b).
Besides reproducing with high accuracy the typical phase diagram of the experimental flow regimes (see Fig.14b), the color gradient LB quantitatively captures the behavior of the droplet diameter, which follows a power law of the flow rate over almost two decades.Cubaud and Mason (2008) (Fig.14c).
Figure 14: (a) Different flow regimes captured by the color gradient LB. From top to bottom: dripping (red aster), jetting (green triangle and white diamond), and tubing (yellow circle). (b) Phase diagram of the different regimes on the $Ca_{1}-Ca_{2}$ plane, where $Ca_{1}$ is the capillary number of the dispersed phase and $Ca_{2}$ of the continuous one. (c) Plot of the normalized droplet diameter $D/H$ versus the flow ratio $Q_{c}/Q_{d}$. Numerical results (red triangles) and experimental ones from Ref.Cubaud and Mason (2008) follow a power law of the flow rate with exponent $\simeq 0.5$. The figure is adapted, with permission, from Ref.Montessori et al. (2019c).
Flow focuser.
A more complex geometry is that of a flow-focuser, where an additional orifice separating the inlet and the outlet chambers causes the droplet generation. Marmottant and Raven (2009); Raven and Marmottant (2009); Garstecki et al. (2004). A sketch of the device used in our simulations and the droplet arrangement observed in the channel are shown in Fig.3, where the resulting emulsion is stabilized by dispersing surfactant molecules preventing coalescence. These devices are particularly suited for the design of highly monodisperse droplets which are of relevance, for example, for the assembly of droplet-based soft templates with a well-defined structure Nan et al. (2024); Xu et al. (2005); Costantini et al. (2014).
@@
A typical example are soft flowing crystals (also termed as microfluidic crystals Marmottant and Raven (2009)), where monodisperse fluid droplets added in a self-repeating way are found to self-assemble in regular patterns, closely resembling the hexagonal order of solids. In LB simulations, droplet formation can be controlled by tuning the ratio of dispersed-to-continuous inlet flow, defined as $\chi=u_{d}/2u_{c}$ (where $u_{d}$ and $u_{c}$ are the speed of dispersed and continuous phases) and the dimensionless near-contact number ${\cal N}_{c}=A\Delta x/\sigma$ (see also Eq.42). Here, $A$ is a constant setting the magnitude of the repulsive force and $\Delta x$ is the minimum distance between two interfaces in close contact, generally ranging from $10$ to $100$ nanometers. If ${\cal N}_{c}\ll 1$ droplets coalesce, otherwise repulsive near-contact forces prevail and droplet merging is inhibited.
@@
In Fig.15a-d, we show a variety of crystal-like structures confined in microfluidic channels obtained from three-dimensional LB simulations Montessori et al. (2019b, 2020) and experiments Marmottant and Raven (2009); Raven and Marmottant (2009) of air bubbles in water. Despite the difference in density ratio, the two systems exhibit very similar features. In particular, the emulsion displays a number of phases comprising a three-row configuration (hex-three, (a)), where a regular array of droplets accommodates along three parallel rows, two two-row ones (wet (b) and dry (c) hex-two) and a single-row structure (hex-one, (d)), in which highly-packed droplets arrange in a neat line. Besides being capable of describing the formation and ordering properties of fluid emulsions under confinement, LB simulations also capture the behavior of the experimental flow rate curve and, remarkably, the non-trivial transition from the hex-two to the hex-one regime (see Fig.15e,f). Indeed, experiments show that, for a gas-liquid foam, $Q_{g}\propto P_{g}^{\alpha}$, where $Q_{g}$ and $P_{g}$ are flow rate and pressure of the gas, and $\alpha=3/2$ for small values of $P_{g}$ Marmottant and Raven (2009). On the contrary, for higher values of $P_{g}$, $Q_{g}$ displays a sharp decrease due to the augmented friction of the interfaces with the walls, an effect signaling the transition to the hex-one phase (Fig.15e). These results are reproduced, with high accuracy, by LB simulations of fluid emulsions, where the flow rate $Q_{d}$ of the dispersed phase follows a power law of $\chi$ (which plays the same role as the gas pressure) with an almost equal value of the exponent. In agreement with experiments, simulations also show that the transition from the hex-two to the hex-one phase is marked by a T1 event Graner et al. (2008); Marmottant et al. (2008); Weaire and Hutzler (2001) (corresponding to a topological rearrangement of neighboring droplets) and a transition front, separating the two phases and comoving with the flowing emulsion.
@@
Figure 15: (a)-(d) Simulated (top rows) and experimental (bottom rows) structures of flowing droplets in a microfluidic channel for $\chi=1$ (a) (hex-three), $\chi=1.5$ (b) (wet hex-two), $\chi=2$ (c) (dry hex-two) and $\chi=3.6$ (d) (hex-one). Here $\chi=u_{d}/2u_{c}$, being $u_{d}$ and $u_{c}$ the speed of dispersed and continuous phases. (e)-(f) Flow curves of experimental Marmottant and Raven (2009) and simulated Montessori et al. (2020) soft flowing crystals. In (e) $P_{g}$ and $Q_{g}$ are the gas pressure and gas flow rate. For values of $P_{g}$ lower than $4$kPa, the mixture is in the hex-two regime, where the flow rate scales as $Q_{g}\propto P^{\alpha}$ and $\alpha\simeq 3/2$. At $P_{g}\simeq 4$kPa, $Q_{g}$ displays a sharp transition towards the hex-one regime. In (f) $Q_{d}=\nu V_{d}$ is the flow rate of the dispersed phase, where $\nu$ is the frequency of droplet generation and $V_{d}$ is the volume of the generated droplet. The flow rate follows a power law with exponent $\simeq 3/2$ for low values of $\chi$, while for $\chi\simeq 3.5$ a decrease of $Q_{d}$ marks the transition from the hex-two to the hex-one phase. The transition front is indicated by a white dotted line in the top inset. The simulations are carried out on a 3D box of $420\times 80\times 20$ lattice nodes. The figures are reproduced with permission from Ref.Montessori et al. (2019b, 2020).
Divergent channel.
@@
LB simulations of microflows have also been capable of describing morphology and assembly of droplets flowing in convergent and divergent channels Montessori et al. (2019b); Gai et al. (2022); Y. et al. (2016); Luo et al. (2021); Montessori et al. (2021a, b), as the ones shown in Fig.16, albeit in these cases mostly on a qualitative ground. More specifically, numerical results suggest that the degree of monodispersity and the arrangement of a confined dense emulsion can be controlled by properly tuning the opening angle $\theta$ of a divergent channel and Capillary number $Ca$ (see Fig.16). Indeed, while for $\theta\lesssim 40^{\circ}$ the emulsion is generally monodisperse (Fig.16a) regardless of the values of $Ca$, for $45^{\circ}\lesssim\theta\lesssim 60^{\circ}$ the generated emulsion becomes bidisperse, an effect amplified at high values of $Ca$ (see Fig.16b,c) because breakup events are more likely to occur. The mechanics leading to droplet rupture is highlighted in Fig.16e-n. The orange droplet gradually stretches because of the combined effect of i) the normal force of the incoming droplet (yellow), ii) the extensional force of the channel expansion, and iii) the downstream droplet (red) acting as a ”wall”. Once a critical elongation is overcome, the orange droplet fragments produce two smaller droplets moving towards the upper and lower part of the channel. Then the process self-repeats; the yellow droplet fills the gap left by the orange one and acts as a ”wall” droplet, while a newly generated one will stretch and finally break following a protocol similar to the orange one. Interestingly, further increasing $\alpha$ restores a monodisperse emulsion, once again regardless of the values of $Ca$ (see Fig.16d). This counter-intuitive behavior is due to the fast recovery of the circular shape in the main channel because of larger aperture angles. Such an expansion favors the transversal displacement of the incoming droplet, thus suppressing the squeezing effect. Note finally that, for all geometries considered, increasing $Ca$ generally leads to the formation of a foamy-like emulsion; this is because, in this regime, morphological deformations become easier and, once breakups occur, the resulting smaller droplets fill the voids among neighboring ones.
@@
Figure 16: (a)-(d). Droplet assembly in diverging channels of opening angle $\theta$ equal to $22.5^{\circ}$ (a), $45^{\circ}$ (b), $60^{\circ}$ (c), $90^{\circ}$ (d) and three different values of $Ca$. The histograms show the distribution of droplet diameters $D_{e}$ for $Ca=0.04$ (dashed line), $Ca=0.1$ (dotted line) $Ca=0.16$ (full line). Note that, for $45^{\circ}\lesssim\theta\lesssim 60^{\circ}$, the emulsion displays a bidisperse structure (disordered), while for values out of that range, the emulsion is essentially monodisperse (ordered). (e)-(n). Pinch-off mechanism from LB simulations (e-h) and experiments (i-n) at $Ca\simeq 0.08$ and $\theta=45^{\circ}$. The split of the orange droplet is caused by the confining effect of the yellow droplet (the ”pincher”) and the red one (the ”wall”). The figure is reproduced with permission from Ref.Montessori et al. (2021b).
@@ V.2 Soft granular media
Figure 17: Left panel, (a)-(c). Phase diagram of a soft granular medium, where $N$ is the number of cores and $\phi_{d}$ is the volume fraction they occupy. Insets (a), (b) and (c) represent three possible configurations, where (a) indicates weakly deformed shells, (b) highly deformed ones, and (c) a linear chain. Right panel, (d)-(e). Experimental realization of a soft granular medium (d) formulated using three Newtonian fluids. A result obtained from LB simulations (e) using three-component fluid, i.e. the dispersed phase (white droplets), the lubricating phase (blue), and an external fluid (black). Copyright granted under the CC 4.0 International license and CC 3.0 license respectively. No changes were made from the original figure from Ref.Montessori et al. (2021c); Guzowski and Garstecki (2015).
[…]
Such a picture decisively depends on the grain structure as well as on the deformability of the cores, features controlling the formation of highly packed structures with a volume fraction typically well above the close packing limit of hard spheres.
@@
Figure 18: (a) Dynamic patterns observed in a soft granular medium by varying $Q_{c}/Q_{d}$. The medium consists of a water-in-oil emulsion surrounded by a further immiscible phase (fluorinated fluid). The left column shows the simulation results (on a 3d lattice of size $280\times 140\times 30$), the central one shows the corresponding experiments and the right one refers to a water (blue)-oil (transparent) mixture. The color bar refers to the fluid velocity in simulations. (b) Mechanism leading to the jet breakup. If a doublet enters the narrowing and unfolds, the droplet chain is stabilized. Otherwise, high-velocity gradients in the surrounding of the doubled (see the pink arrow) destabilize the chain. Copyright granted under the CC 4.0 International license. No changes were made from the original figure from Ref.Bogdan et al. (2022).
LB simulations prove, once again, instrumental to capture crucial aspects of the physics of such systems. In Fig.18a, for example, a series of experiments about the formation of dynamic modes of a confined granular medium (a densely packed monodisperse emulsion focused by an external immiscible phase through an orifice) with $\phi_{d}\simeq 0.8$ are presented, while three dimensional numerical simulations are performed using the color gradient LB described in section III.3.1. Upon varying the flow ratio $Q_{c}/Q_{d}$ between the continuous and the dispersed phases, both experiments and simulations show that one can distinguish four modes, consisting of (i) a thick, weakly oscillating jetting phase for $Q_{c}/Q_{d}\simeq 1$, (ii) a thin, strongly oscillating jet with occasional breakups for $Q_{c}/Q_{d}\simeq 2$, (iii) an irregular dripping phase of polydisperse double emulsions for $Q_{c}/Q_{d}\simeq 8$, and (iv) a regular dripping phase of monodisperse double emulsions for $Q_{c}/Q_{d}\simeq 15$. It is worth highlighting that the experimental values of the flow ratio match almost one-to-one the numerical values (see Fig.18a), with differences very likely caused by slightly different geometrical parameters (such as length and width of the constriction) and volume fractions of the cores. In stark contrast, Newtonian liquids (such as water in oil mixture) only show two primary modes, i.e. dripping and jetting, although other regimes have been reported in experiments Utada et al. (2007).
Besides describing at a high level of accuracy these complex structures, LB simulations also unveil the mechanism causing the jet breakup and leading to the dripping phase, as shown in Fig.18b. This is essentially based on the droplet rearrangement upon entering the constriction and on the magnitude of the associated velocity gradients. Indeed, when a doublet (i.e. a pair of droplets) enters the narrowing, the surrounding continuous phase accelerates to conserve the flux. If the doublet unfolds into a chain, the velocity gradients remain weak and the chain is stabilized; on the contrary, if the doublet survives, increasing velocity gradients around it destabilize the jet leading to the breakup.